OpenStudy (habbababba1):

If f(x) = 3x + 2 and g(x) = 2x – 2, what is (f – g)(x)? x + 4 x – 2 x 5x – 2 x – 4

OpenStudy (habbababba1):

will fan and medal!!

OpenStudy (habbababba1):

@Error1603

OpenStudy (habbababba1):

@Ebonygotthejuice

OpenStudy (ebonygotthejuice):

idk

OpenStudy (habbababba1):

oh well atleast u responded

OpenStudy (ebonygotthejuice):

lol i mean i'd just guess

OpenStudy (habbababba1):

i have to have every answer right to go on to the next test

OpenStudy (ebonygotthejuice):

im sorry

OpenStudy (habbababba1):

its ok ig haha nothin can do but wait

OpenStudy (habbababba1):

@ONYXxELITE

OpenStudy (onyxxelite):

Yes @habbababba1

OpenStudy (habbababba1):

can u help with this apperantly no one else knows how

OpenStudy (onyxxelite):

hold on let me read it

OpenStudy (habbababba1):

ok

OpenStudy (onyxxelite):

g(f(x)) means that the function g is composed at the function f. You need to plug the value of f(x) into the function g and simplify. g(x) = 2x - 1 Replace x with the value of f(x) which is 3x - 2 and simplify. g(f(x)) = 2x - 1 g(3x - 2) = 2(3x - 2) - 1 g(3x - 2) = 6x - 4 - 1 g(3x - 2) = 6x - 5 So, g(f(x)) = 6x - 5